### v1

@@ Abstract
For a relational structure ${\mathbb{X}}$ we investigate the partial order $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$, where ${\mathbb{P}}({\mathbb{X}}):=\{f[X]:f\in\mathop{\rm Emb}\nolimits({\mathbb{X}})\}$. A previous analysis shows that ${\mathfrak{h}}=\omega_{1}$ implies that for each countable ordinal we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin})$. But in ZFC we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))\cong\prod_{i=0}^{n}((\mathop{\rm rp}\nolimits^{r_{i}}(P(\omega^{\gamma_{i}})/{\mathcal{I}}_{\omega^{\gamma_{i}}}))^{+})^{s_{i}}$, where $\sum_{i=n}^{0}\omega^{\gamma_{i}+r_{i}}s_{i}+k$ is the presentation of $\alpha$ in the Cantor normal form and $\mathop{\rm rp}\nolimits({\mathbb{B}})$ denotes the reduced power of a Boolean algebra ${\mathbb{B}}$ modulo finite: ${\mathbb{B}}^{\omega}/\mathop{\rm Fin}$. Consequently, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin})^{+}\ast\pi)$, where $\pi$ is an $P(\omega)/\mathop{\rm Fin}$-name for an $\sigma$-closed separative atomless poset.
Here we consider uncountable ordinals. Since $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is isomorphic to the direct product $\prod_{i=1}^{n}(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta_{i}}))^{s_{i}}$, where $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is the Cantor normal form for $\alpha$, the analysis is reduced to the investigation of the posets of the form ${\mathbb{P}}(\omega^{\delta})$. It turns out that, in ZFC, either the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed and completely embeds $P(\omega)/\mathop{\rm Fin}$ and, hence, preserves $\omega_{1}$ and forces $|{\mathfrak{c}}|=|{\mathfrak{h}}|$, or, otherwise, completely embeds the algebra $P(\lambda)/[\lambda]^{<\lambda}$, for some regular $\omega<\lambda\leq\mathop{\rm cf}\nolimits(\delta)$, and collapses $\omega_{2}$ to $\omega$. Regarding the Cantor normal form, the first case appears iff for each $i\leq n$ we have $\mathop{\rm cf}\nolimits(\delta_{i})\leq\omega$, or $\delta_{i}=\theta_{i}+\mathop{\rm cf}\nolimits(\delta_{i})$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta_{i}\geq\mathop{\rm cf}\nolimits(\delta_{i})>\mathop{\rm cf}\nolimits(\theta_{i})=\omega$ and $\theta_{i}=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\mathop{\rm cf}\nolimits(\delta_{i})$, for all $n\in\omega$.
@@ 1 Introduction
In this paper we continue the investigation of posets of copies of ordinals initiated in [14], where countable ordinals were considered. Generally speaking, if ${\mathbb{X}}$ is a relational structure and $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is the set of its self-embeddings, $f:{\mathbb{X}}\hookrightarrow{\mathbb{X}}$, then the partial order $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$, where ${\mathbb{P}}({\mathbb{X}}):=\{f[X]:f\in\mathop{\rm Emb}\nolimits({\mathbb{X}})\}$, is called the poset of copies of ${\mathbb{X}}$. Denoting by $\mathop{\rm sq}\nolimits({\mathbb{P}})$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))$ the separative quotient and the Boolean completion of a partial order ${\mathbb{P}}$, the implications
@@
express the relationship between different “similarity relations” in the class of relational structures and each of them provides a classification of structures [12, 15]. The last one, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})$, gives the coarsest classification of structures, it is equivalent to the forcing equivalence of posets of copies, ${\mathbb{P}}({\mathbb{X}})\equiv_{forc}{\mathbb{P}}({\mathbb{Y}})$, and, hence, it can be explored using forcing-theoretic methods and known facts from set theory. The related results can be regarded as the statements belonging to the theory of relations, set theory or theory of Boolean algebras.
It turns out that, when the posets of copies of countable linear orders are in question, the main roles are played by the poset ${\mathbb{S}}$ of perfect subsets of the real line ordered by inclusion (the Sacks forcing) and the algebra $P(\omega)/\mathop{\rm Fin}\nolimits$. Namely, if ${\mathbb{X}}$ is a non-scattered linear order (i.e, if ${\mathbb{Q}}\hookrightarrow{\mathbb{X}}$), then $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}}\ast\pi)$, where $\pi$ is an ${\mathbb{S}}$-name for a $\sigma$-closed poset and, consistently, the name for $P(\omega)/\mathop{\rm Fin}\nolimits$ in the Sacks extension [17]. On the other hand, if ${\mathbb{X}}$ is a scattered linear order, then the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})$ is $\sigma$-closed, atomless and, consistently, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ [13]. Well orders were investigated in [14] and the main results will be presented later in this section. We note that for the real line ${\mathbb{R}}$ we have $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{R}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}})$ [16] and that the Sacks forcing appears as a factor in the presentation of posets of copies of several ultrahomogeneous structures, [8, 17, 18, 19, 20, 21].
@@
The results obtained in this area can be applied in investigation of transformation monoids. If ${\mathbb{X}}$ is a relational structure and $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),\circ,\mathop{\rm id}\nolimits_{X}\rangle$ is the monoid of its self-embeddings, first we note that the set $\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})$ of automorphisms of ${\mathbb{X}}$ together with a choice of an embedding $f_{A}\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$ such that $f_{A}[X]=A$, for each $A\in{\mathbb{P}}({\mathbb{X}})$, completely determine the set of embeddings: $\mathop{\rm Emb}\nolimits({\mathbb{X}})=\{f_{A}\circ a:A\in{\mathbb{P}}({\mathbb{X}})\mbox{ and }a\in\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})\}$. Second, we recall that right Green’s preorder on $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is defined by $f\preccurlyeq^{R}g$ iff $f\circ h=g$, for some $h\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$, and note that, defining $f\thickapprox^{R}g$ iff $f\preccurlyeq^{R}g$ and $g\preccurlyeq^{R}f$, we have
$f\preccurlyeq^{R}g\Leftrightarrow g[X]\subset f[X]\;\;\mbox{ and }\;\;f\thickapprox^{R}g\Leftrightarrow g[X]=f[X].$ (1)
Thus the poset of copies $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$ is exactly the antisymmetric quotient of the preorder $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),(\preccurlyeq^{R})^{-1}\rangle$.
Now we turn to the particular case considered in this paper, when ${\mathbb{X}}$ is a well order, say an ordinal $\alpha$. Then it is a rigid structure ($\mathop{\mathrm{Aut}}\nolimits(\alpha)=\{\mathop{\rm id}\nolimits_{\alpha}\}$) and for each copy $A\in{\mathbb{P}}(\alpha)$ there is a unique $f_{A}\in\mathop{\rm Emb}\nolimits(\alpha)$ such that $f_{A}[\alpha]=A$. Thus we have a simple reconstruction:
$\mathop{\rm Emb}\nolimits(\alpha)=\{f_{A}:A\in{\mathbb{P}}(\alpha)\}.$
Moreover, defining the binary operation $\ast$ on ${\mathbb{P}}(\alpha)$ by $A\ast B=f_{A}[B]$ and the mapping $F:{\mathbb{P}}(\alpha)\rightarrow\mathop{\rm Emb}\nolimits(\alpha)$ by $F(A)=f_{A}$, we obtain an isomorphism of monoids:
$\langle{\mathbb{P}}(\alpha),\ast,\alpha\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\circ,\mathop{\rm id}\nolimits_{\alpha}\rangle.$ (2)
In addition, the preorder $\preccurlyeq^{R}$ on $\mathop{\rm Emb}\nolimits(\alpha)$ becomes a partial order and the same mapping $F$ witnesses that
$\langle{\mathbb{P}}(\alpha),\supset\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\preccurlyeq^{R}\rangle.$ (3)
For convenience we consider the poset $\langle{\mathbb{P}}(\alpha),\subset\rangle$, shortly denoted by ${\mathbb{P}}(\alpha)$.
@@
In [14] the posets of copies of countable ordinals $\alpha\geq\omega$ were investigated and, in order to present the main results, we explain some notation. If ${\mathbb{B}}$ is Boolean lattice, $\mathop{\rm rp}\nolimits({\mathbb{B}})$ denotes the reduced power ${\mathbb{B}}^{\omega}/\mathop{\rm Fin}\nolimits$, and the iterated reduced powers $\mathop{\rm rp}\nolimits^{n}({\mathbb{B}})$, for $n\in\omega$, are defined by $\mathop{\rm rp}\nolimits^{0}({\mathbb{B}})={\mathbb{B}}$ and $\mathop{\rm rp}\nolimits^{n+1}({\mathbb{B}})=\mathop{\rm rp}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))$. If $\delta>0$ is an ordinal, then ${\mathcal{I}}_{\omega^{\delta}}:=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}\subset P(\omega^{\delta})$ is an ideal and the Boolean algebra $P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$ is the corresponding quotient. The first result shows that the separative quotient of the poset ${\mathbb{P}}(\alpha)$ is isomorphic to the forcing (free) product of finitely many Boolean algebras.
@@
If $\alpha$ is a countable ordinal presented in the Cantor normal form,
$\alpha=\omega^{\gamma_{n}+r_{n}}s_{n}+\dots+\omega^{\gamma_{0}+r_{0}}s_{0}+k,$ (4)
where $k\in\omega$, $r_{i}\in\omega$, $s_{i}\in{\mathbb{N}}$, $\gamma_{i}\in\mathop{\rm Lim}\nolimits\cup\{1\}$ and $\gamma_{n}+r_{n}>\dots>\gamma_{0}+r_{0}$, then
So, inductively, for $\alpha=\omega$ from (5) we obtain $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$; further we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega s)=((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})^{s}$, for $s\in{\mathbb{N}}$; $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega\omega)=(\mathop{\rm rp}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits))^{+}$; and, generally, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{r}))=(\mathop{\rm rp}\nolimits^{r-1}(P(\omega)/\mathop{\rm Fin}\nolimits))^{+}$, for $r\in{\mathbb{N}}$. Further for a limit $\gamma$ and $r\in\omega$ we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\gamma+r}))=(\mathop{\rm rp}\nolimits^{r}(P(\omega^{\gamma})/{\mathcal{I}}_{\omega^{\gamma}}))^{+}$; and, finally, by (5) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is the free product of the posets $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha_{i}))$, where $\alpha_{i}=\omega^{\gamma_{i}+r_{i}}$, for $i\leq n$, are the infinite summands from (4).
The second result is related to the aforementioned coarse classification of structures (corresponding to the similarity $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\beta)))$) and, by (3), a classification of their embedding monoids. We recall that the distributivity number ${\mathfrak{h}}({\mathbb{P}})$ of a partial order ${\mathbb{P}}$ is the minimal size of a family ${\mathcal{D}}$ of open dense subsets of ${\mathbb{P}}$ such that $\bigcap{\mathcal{D}}$ is not dense, and that $\omega_{1}\leq{\mathfrak{h}}:={\mathfrak{h}}((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})\leq{\mathfrak{c}}$.
Theorem 1.2
If $\alpha\geq\omega+\omega$ is a countable ordinal, then we have
  • (a)
    ${\mathbb{P}}(\alpha)\equiv_{forc}(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi$, where $[\omega]\Vdash$$\pi$ is an $\sigma$-closed separative atomless poset”; and, consequently,
    $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi);$ (6)
  • (b)
    If ${\mathfrak{h}}=\omega_{1}$, then ${\mathbb{P}}(\alpha)\equiv_{forc}(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$ and, hence,
    $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+}).$ (7)
@@
The following example shows that the statement that (7) holds for all countable ordinals is independent of ZFC. So the coarse classification depends on the model of ZFC in which we live. Defining ${\mathfrak{h}}_{n}:={\mathfrak{h}}(((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})^{n})$, for $n\in{\mathbb{N}}$, we have ${\mathfrak{h}}_{1}\geq{\mathfrak{h}}_{2}\geq{\mathfrak{h}}_{3}\geq\dots\geq\omega_{1}$ and, by Theorem 1.1, ${\mathfrak{h}}(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega n))={\mathfrak{h}}_{n}$. Shelah and Spinas proved in [24] that for each $n\in{\mathbb{N}}$ there is a model of ZFC in which ${\mathfrak{h}}_{n+1}<{\mathfrak{h}}_{n}$, which implies that ${\mathbb{P}}(\omega n)\not\equiv_{forc}{\mathbb{P}}(\omega(n+1))$ and, hence, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega n)))\not\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega(n+1))))$.
In this paper we investigate the posets ${\mathbb{P}}(\alpha)$ for uncountable ordinals $\alpha$. Section 2 contains definitions and known facts used in the paper.
@@
In Section 3 we show that the separative quotient $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is isomorphic to the direct product of $\sum_{i=1}^{n}s_{i}$-many factors $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta_{i}})$, where $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is the Cantor normal form for $\alpha$, and reduce the analysis to the investigation of these factors.
@@
So, in Section 4 we regard the posets of the form ${\mathbb{P}}(\omega^{\delta})$, for ordinals $\delta>0$. It turns out that the behavior of ${\mathbb{P}}(\omega^{\delta})$ depends on the cofinality of $\delta$. If $\mathop{\rm cf}\nolimits(\delta)\in\{1,\omega\}$ (Case (A)), then the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed, completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$ and, hence, collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$. Otherwise we have $\mathop{\rm cf}\nolimits(\delta):=\kappa\geq\omega_{1}$ and, presenting $\delta$ in its Cantor normal form in the base $\kappa$, we detect additional four cases, (B)–(E). If (B) holds, then again, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. In Cases (D) and (E) the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds ${\mathcal{P}}_{\kappa}^{+}$, where ${\mathcal{P}}_{\kappa}$ denotes the algebra $P(\kappa)/[\kappa]^{<\kappa}$, and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ (and, hence, $\omega_{2}$) to $\omega$. Finally, in Case (C) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds ${\mathcal{P}}_{\lambda}^{+}$, where $\omega_{1}<\lambda<\kappa$, and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})$ (and, again, $\omega_{2}$) to $\omega$.

### v2

@@ Abstract
The set of copies ${\mathbb{P}}(\alpha):=\{f[\alpha]:f\in\mathop{\rm Emb}\nolimits(\alpha)\}$ of an ordinal $\alpha$ reconstructs the monoid $\mathop{\rm Emb}\nolimits(\alpha)$ of self embeddings of $\alpha$ and $\langle{\mathbb{P}}(\alpha),\subset\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\succcurlyeq^{R}\rangle$, where $\preccurlyeq^{R}$ is right Green’s preorder. Regarding the classification of countable ordinals up to isomorphism of the Boolean completions of their posets of copies we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin})^{+}\ast\pi)$, where $\pi$ is a $P(\omega)/\mathop{\rm Fin}$-name for a $\sigma$-closed separative poset; in particular, under ${\mathfrak{h}}=\omega_{1}$ we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin})$, whenever $\alpha<\omega_{1}$.
Here we consider ordinals $\alpha\geq\omega_{1}$ and show that the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is either $\sigma$-closed and completely embeds a finite power of $(P(\omega)/\mathop{\rm Fin})^{+}$, or completely embeds $(P(\kappa)/[\kappa]^{<\kappa})^{+}$, for some regular cardinal $\kappa>\omega$ and collapses $\omega_{2}$ to $\omega$. It is consistent that the algebra $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))$ is isomorphic to $\mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}})$ or to $\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$, for each infinite ordinal $\alpha<{\mathfrak{c}}=\omega_{2}$. More generally, the partial order $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is isomorphic to a forcing product of the algebras of the form $P(\omega^{\delta})/{\mathcal{I}}_{\delta}$ and we detect several classes of ordinals $\delta$ such that the algebra $\mathop{\rm ro}\nolimits(P(\omega^{\delta})/{\mathcal{I}}_{\delta})$ is isomorphic to a collapsing algebra.
@@ 1 Introduction
In this paper we continue the investigation of the posets of copies of ordinals initiated in [10]. Generally speaking, if ${\mathbb{X}}$ is a relational structure and $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is the set of its self-embeddings, then the partial order $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$, where ${\mathbb{P}}({\mathbb{X}}):=\{f[X]:f\in\mathop{\rm Emb}\nolimits({\mathbb{X}})\}$, is called the poset of copies of ${\mathbb{X}}$. Denoting by $\mathop{\rm sq}\nolimits({\mathbb{P}})$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))$ the separative quotient and the Boolean completion of a partial order ${\mathbb{P}}$, the implications
@@
express the relationship between “similarity relations” in the class of relational structures and each of them provides a classification of structures [8, 11]. The last one, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})$, gives the coarsest classification, it is equivalent to the forcing equivalence of posets of copies, and, hence, it can be explored using forcing-theoretic methods. The related results can be regarded as the statements belonging to the theory of relations, set theory or the theory of Boolean algebras.
@@
In addition, these results can be applied in investigation of the transformation monoids $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),\circ,\mathop{\rm id}\nolimits_{X}\rangle$. First we note that the set $\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})$ of automorphisms of ${\mathbb{X}}$ together with a choice of an embedding $f_{A}\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$ such that $f_{A}[X]=A$, for each $A\in{\mathbb{P}}({\mathbb{X}})$, determine the set of embeddings: $\mathop{\rm Emb}\nolimits({\mathbb{X}})=\{f_{A}\circ a:A\in{\mathbb{P}}({\mathbb{X}})\mbox{ and }a\in\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})\}$. Second, we recall that right Green’s preorder on $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is defined by $f\preccurlyeq^{R}g$ iff $f\circ h=g$, for some $h\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$, and note that, defining $f\thickapprox^{R}g$ iff $f\preccurlyeq^{R}g$ and $g\preccurlyeq^{R}f$, we have $f\preccurlyeq^{R}g\Leftrightarrow g[X]\subset f[X]$ and $f\thickapprox^{R}g\Leftrightarrow g[X]=f[X]$. So, $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$ is the antisymmetric quotient of the preorder $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),\succcurlyeq^{R}\rangle$. If, in particular, ${\mathbb{X}}$ is an ordinal $\alpha$, it is a rigid structure and for each copy $A\in{\mathbb{P}}(\alpha)$ there is a unique $f_{A}\in\mathop{\rm Emb}\nolimits(\alpha)$ such that $f_{A}[\alpha]=A$. Thus, here we have $\mathop{\rm Emb}\nolimits(\alpha)=\{f_{A}:A\in{\mathbb{P}}(\alpha)\}$ and defining the binary operation $\ast$ on ${\mathbb{P}}(\alpha)$ by $A\ast B=f_{A}[B]$ and the mapping $F:{\mathbb{P}}(\alpha)\rightarrow\mathop{\rm Emb}\nolimits(\alpha)$ by $F(A)=f_{A}$, we obtain an isomorphism of monoids; namely, $F:\langle{\mathbb{P}}(\alpha),\ast,\alpha\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\circ,\mathop{\rm id}\nolimits_{\alpha}\rangle$. In addition, the preorder $\preccurlyeq^{R}$ on $\mathop{\rm Emb}\nolimits(\alpha)$ becomes a partial order and the same mapping $F$ witnesses that $\langle{\mathbb{P}}(\alpha),\subset\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\succcurlyeq^{R}\rangle$.
It turns out that, regarding countable linear orders, the main roles in the description of the algebra $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})$ are played by the poset ${\mathbb{S}}$ of perfect subsets of the real line ordered by inclusion (the Sacks forcing) and the algebra $P(\omega)/\mathop{\rm Fin}\nolimits$. Namely, if ${\mathbb{X}}$ is non-scattered (i.e., ${\mathbb{Q}}\hookrightarrow{\mathbb{X}}$), then $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}}\ast\pi)$, where $\pi$ is an ${\mathbb{S}}$-name for a $\sigma$-closed poset and, consistently, the name for $P(\omega)/\mathop{\rm Fin}\nolimits$ in the Sacks extension [14]. On the other hand, if ${\mathbb{X}}$ is a scattered linear order, then the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})$ is $\sigma$-closed, atomless and, consistently, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ [9]. We note that for the real line ${\mathbb{R}}$ we have $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{R}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}})$ [12] and that the Sacks forcing appears as a factor in the presentation of posets of copies of several ultrahomogeneous structures, [5, 14, 15, 16, 17, 18].
@@
In [10] the posets of copies of countable ordinals $\alpha\geq\omega$ were investigated and, in order to present the main results, we explain some notation. If ${\mathbb{B}}$ is Boolean lattice, $\mathop{\rm rp}\nolimits({\mathbb{B}})$ denotes the reduced power ${\mathbb{B}}^{\omega}/\mathop{\rm Fin}\nolimits$, and the iterated reduced powers $\mathop{\rm rp}\nolimits^{n}({\mathbb{B}})$, for $n\in\omega$, are defined by $\mathop{\rm rp}\nolimits^{0}({\mathbb{B}})={\mathbb{B}}$ and $\mathop{\rm rp}\nolimits^{n+1}({\mathbb{B}})=\mathop{\rm rp}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))$. If $\delta>0$ is an ordinal, then ${\mathcal{I}}_{\omega^{\delta}}:=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}\subset P(\omega^{\delta})$ is an ideal and the Boolean algebra $P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$ is the corresponding quotient. The first result shows that the separative quotient of the poset ${\mathbb{P}}(\alpha)$ is isomorphic to the forcing product of finitely many Boolean algebras: if $\alpha=\omega^{\gamma_{n}+r_{n}}s_{n}+\dots+\omega^{\gamma_{0}+r_{0}}s_{0}+k$ is a countable ordinal presented in the Cantor normal form, where $k\in\omega$, $r_{i}\in\omega$, $s_{i}\in{\mathbb{N}}$, $\gamma_{i}\in\mathop{\rm Lim}\nolimits\cup\{1\}$ and $\gamma_{n}+r_{n}>\dots>\gamma_{0}+r_{0}$, then
The second result is related to the aforementioned coarse classification of ordinals and their embedding monoids. We recall that the distributivity number ${\mathfrak{h}}({\mathbb{P}})$ of a partial order ${\mathbb{P}}$ is the minimal size of a family ${\mathcal{D}}$ of open dense subsets of ${\mathbb{P}}$ such that $\bigcap{\mathcal{D}}$ is not dense, and that $\omega_{1}\leq{\mathfrak{h}}:={\mathfrak{h}}((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})\leq{\mathfrak{c}}$.
@@
If $\alpha$ is a countable ordinal, then
(a) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi)$, where $[\omega]\Vdash$$\pi$ is a $\sigma$-closed separative poset”;
(b) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})$, if $\,{\mathfrak{h}}=\omega_{1}$.
@@
In this paper we investigate the posets ${\mathbb{P}}(\alpha)$ for uncountable ordinals $\alpha$. Section 2 contains definitions and known facts used in the paper. In Section 3 we show that the separative quotient $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is isomorphic to the direct product of $\sum_{i=1}^{n}s_{i}$-many factors $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta_{i}})$, where $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is the Cantor normal form for $\alpha$, and reduce the analysis to the investigation of these factors.
@@
So, in Section 4 we regard the posets of the form ${\mathbb{P}}(\omega^{\delta})$. It turns out that the behavior of ${\mathbb{P}}(\omega^{\delta})$ depends on the cofinality $\kappa:=\mathop{\rm cf}\nolimits(\delta)$. If $\kappa\leq\omega$ (Case (A)), then the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed, completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$ and, hence, collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$. Otherwise we have $\kappa\geq\omega_{1}$ and, presenting $\delta$ in its Cantor normal form in the base $\kappa$, we detect additional four cases, (B)–(E). If (B) holds, then again, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. In Cases (D) and (E) the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds the poset ${\mathcal{P}}_{\kappa}:=(P(\kappa)/[\kappa]^{<\kappa})^{+}$ and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ (and, hence, $\omega_{2}$) to $\omega$. Finally, in Case (C) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds ${\mathcal{P}}_{\lambda}$, where $\omega<\lambda<\kappa$, and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})$ (and, again, $\omega_{2}$) to $\omega$.